# Simulating quantum computation: how many “bits” for “it”?

### v2

@@ III.2 General case
Now, the size of every Clifford orbit of vertices of $\Lambda_{n}$ is upper-bounded by the size $|\mathcal{C}\ell_{n}|$ of the n-qubit Clifford group. Therefore, the number of vertices reachable by Algorithm 1 is bounded by $2^{2n^{2}+3n}\times|\mathcal{C}\ell_{n}|$. The number of bits required to specify an $n$-qubit Clifford gate is no more than $2n^{2}+3n$ (namely, $2n\times(2n+1)$ bits to specify the conjugation relations of a generating set of Pauli operators including the sign, minus $2n^{2}-n$ constraints for preserving the Pauli commutation relations). Hence, the number of bits needed to specify any vertex reached in the simulation using to Algorithm 1 is no more than $4n^{2}+6n$, as claimed. $\Box$
IV Discussion
@@ Appendix A Upper bound on the number of vertices of $\Lambda$
The proof relies on the following lemma bounding the number of stabilizer states for $n$ qubits.
Lemma 2.
For any number of qubits $n\in\mathbb{N}$, the number of pure $n$-qubit stabilizer states is $|\mathcal{S}_{n}|\leq 3\cdot 2^{(n^{2}+3n)/2}$.
Proof of Lemma 2. This can be shown by induction. For the base case, we have $|S_{1}|=6\leq 3\cdot 2^{(1^{2}+3)/2}=12$. Now, assume $|\mathcal{S}_{n}|\leq 3\cdot 2^{(n^{2}+3n)/2}$ for some $n\in\mathbb{N}$. Then
$|\mathcal{S}_{n+1}|=$ $1\displaystyle 2^{n+1}\prod\limits_{k=1}^{n+1}2^{k}+$ 1
$=$ $2(2^{n+1}+1)\cdot|\mathcal{S}_{n}|$
$\leq$ $2(2^{n+1}+1)\cdot 3\cdot 2^{(n^{2}+3n)}$
$\leq$ $3\cdot 2^{[(n+1)^{2}+3(n+1)]/2}.$
Therefore, by induction, $|\mathcal{S}_{n}|\leq 3\cdot 2^{(n^{2}+3n)/2}$ for all $n\in\mathbb{N}$. $\Box$
Proof of Lemma 1. According to the upper bound theorem of polytope theory [11] (for a more detailed exposition on the upper bound theorem see Ref. [18, §8.4]), the number of facets of a $D$-dimensional polytope with $v$ vertices is bounded by the number of facets of the $D$-dimensional cyclic polytope with $v$ vertices, denoted $C(v,D)$. By duality, the number of vertices of the $\Lambda$ polytope on $n$ qubits is bounded by the number of vertices of the polar dual of the $14^{n}-$ 1-dimensional cyclic polytope with $|\mathcal{S}_{n}|$ vertices.
The number of vertices of the dual of $C(v,D)$ (equivalently the number of facets of $C(v,D)$) is [24, §4.7],
@@
$\log_{2}(|\mathcal{V}_{n}|)\leq$ $1+2^{2n-1}\left[\log_{2}(|\mathcal{S}_{n}-2^{2n-1})+\log_{2}(e)-2n+1\right]$
$\leq$ $1+2^{2n-1}\left[\log_{2}(3\cdot 2^{(n^{2}+3n)/2})-2n+1+\log_{2}(e)\right]$
@@
This proves the upper bound. $\Box$

### v3

@@ IV.2 General case
Now, the size of every Clifford orbit of vertices of $\Lambda_{n}$ is upper-bounded by the size $|\mathcal{C}\ell_{n}|$ of the n-qubit Clifford group. Therefore, the number of vertices reachable by Algorithm 1 is bounded by $2^{2n^{2}+3n}\times|\mathcal{C}\ell_{n}|$. The number of bits required to specify an $n$-qubit Clifford gate is no more than $2n^{2}+3n$. Namely, the conjugation action of a Clifford unitary on a Pauli operator is described by a binary-valued $2n\times 2n$ matrix plus a $2n$-component binary vector for the signs of Pauli operators [17], requiring $2n\times(2n+1)$ bits. From this, the $2n^{2}-n$ constraints imposed by preservation of the Pauli commutation relations need to be subtracted. Hence, to specify a vertex reachable in the simulation using Algorithm 1, we specify a Clifford orbit $[A_{\nu}]$ and a Clifford operation $g$, which takes no more than $4n^{2}+6n$ bits, as claimed. $\Box$
We may generalize beyond the scenario of Corollary 1 by supplementing the initial magic state |$\ket{H}^{\otimes n}$ with an $m$-qubit ancilla in a stabilizer state, where the stabilizer part can be kept separate and tracked independently in the simulation. As in the previous generalization, the computational power does not increase, but a broader physical setting is represented.
Then the number of bits needed depends on both the number of qubits $m$ supporting the stabilizer part of the input, and the number of magic states $n$, namely,
$[2m^{2}+m]+[4n^{2}+6n]+[2(n+m)^{2}+3(n+m)].$ (4)
Here, the first term accounts for the stabilizer part of the state, the second term for the the magic part (this is the first version of the bound given in the proof of Theorem 2 shown in Fig. 2), and the third term is for the $m+n$ qubit Clifford group element that needs to be tracked, as in the proof of Corollary 1.
In deriving Eq. (4), we make use of the fact that a tensor product of a projector onto an $m$-qubit stabilizer state with a vertex of $\Lambda_{n}$ is a vertex of $\Lambda_{m+n}$ [19, Theorem 2], and we use the stabilizer formalism techniques from Ref. [19, §4] to isolate update of the stabilizer part from the magic part of the state. The number of bits tracked remains quadratic in $n$ and $m$.
V Discussion
@@ Appendix A Upper bound on the number of vertices of $\Lambda$
Proof of Lemma 1. According to the upper bound theorem of polytope theory [12] (for a more detailed exposition on the upper bound theorem see Ref. [20, §8.4]), the number of facets of a $D$-dimensional polytope with $v$ vertices is bounded by the number of facets of the $D$-dimensional cyclic polytope with $v$ vertices, denoted $C(v,D)$. By duality, the number of vertices of the $\Lambda$ polytope on $n$ qubits is bounded by the number of vertices of the polar dual of the $14^{n}-$ 1-dimensional cyclic polytope with $|\mathcal{S}_{n}|$ vertices.
The number of vertices of the dual of $C(v,D)$ (equivalently the number of facets of $C(v,D)$) is [29, §4.7],
@@
$\log_{2}(|\mathcal{V}_{n}|)\leq$ $1+2^{2n-1}\left[\log_{2}\left(|\mathcal{S}_{n}|-2^{2n-1}\right)-2n+1+\log_{2}(e)\right]$
$\leq$ $1+2^{2n-1}\left[\log_{2}\left(5\cdot 2^{(n^{2}+3n)/2}\right)-2n+1+\log_{2}(e)\right]$
@@
Here in the second line we use the upper bound on the number of $n$-qubits stabilizer states, $|\mathcal{S}_{n}|\leq 5\cdot 2^{(n^{2}+3n)/2}$, from Ref. [30]. This proves the upper bound. $\Box$
@@
Appendix B Quantum computation with magic states
Quantum computation with magic states (QCM) is a universal model of quantum computation in which the stabilizer subtheory (stabilizer states, Clifford gates, and Pauli measurements) is supplemented with so-called “magic” (non-stabilizer) states to achieve quantum computational universality [13]. The stabilizer subtheory is not universal for quantum computation, and circuits consisting of elements from this subtheory can be efficiently simulated by a classical computer [24, 31].
In the circuit model, universality is provided by the capacity to perform non-Clifford unitary gates. In the magic state model, on the other hand, universality is restored by the inclusion of non-stabilizer input states. Figure 3 shows a standard example of this phenomenon, a non-Clifford $T$ gate is performed, using only stabilizer operations, by injecting the magic state |$\ket{H}=(\ket{0}+e^{i\pi/4}\ket{1})/\sqrt{2}$. Therefore, given access to operations in the stabilizer subtheory, as well as copies of a suitable magic state such as |$\ket{H}$, any quantum computation can be performed.
For quantum computational universality, it suffices to supplement the stabilizer operations with copies of a single-qudit magic state such as |$\ket{H}$. Of course, without the capacity to perform nonstabilizer operations, it is not clear that these states can be prepared. Fortunately, there exist protocols called magic state distillation protocols in which multiple noisy copies of a magic state are consumed to produce fewer copies of the magic state with higher fidelity. This approach is one of the leading candidates for scalable, fault-tolerant quantum computation [32].
[…]
Figure 3: Quantum circuit showing the implementation of the non-Clifford $T$ gate in the QCM framework, via the injection of the magic state |$\ket{H}=(\ket{0}+e^{i\pi/4}\ket{1})/\sqrt{2}$. Since the Clifford+$T$ gate set is universal for quantum computation, this proves that, with access to suitable magic states, QCM is also universal. Circuit reproduced from [17, Figure 10.25].
For more on quantum computation with magic states, see for example [32, §II] or [17, Ch. 10], or for a more detailed summary see [2, §SM.II].